A nozzle is a device for increasing the velocity of a steadily flowing stream. At the inlet to a certain nozzle, the enthalpy of the fluid passing is 3000 kJ/kg and the velocity is 60 m/s. At the discharge end, the enthalpy is 2762 kJ/kg. The nozzle is horizontal and there is negligible heat loss from it. Calculate:
a. The velocity of the fluid at the exit
1
Expert's answer
2020-02-25T10:22:09-0500

From the conservation of energy:


"h_1+0.5v_1^2=h_2+0.5v_2^2"

"3000\cdot 10^3+0.5(60)^2=2762\cdot 10^3+0.5v_2^2"

"v_2=692.5\frac{m}{s}"